OpenStudy (rsadhvika):

show that 3^a + 3^b + 1 is never a perfect square a, b are natural numbers

OpenStudy (anonymous):

pro is writing

Parth (parthkohli):

Just throwing out stupid ideas. We want to show this is not possible.\[3^a + 3^b = (k+1)(k-1)\]

OpenStudy (anonymous):

If the first term is not a perfect square, the third term is not a perfect square, the middle term is not either 2 or −2 times the product of the square root of the first term and the square root of the third term, then it means it is not a perfect square.

Parth (parthkohli):

Please evaluate it before you copy and paste.

OpenStudy (rsadhvika):

I guess from that we may sthat \(k\) has to be odd

OpenStudy (anonymous):

don't look at me i am chust warming up here

imqwerty (imqwerty):

perfect squares end in either 1/4/9/6/5/0 and we have this- \(3^a+3^b+1\)\[(2+1)^a +(2+1)^b+1\]\(\large\underbrace{\sum_{r=0}^{a} ~^aC_r(2)^r + \sum_{r=0}^{b}~^bC_r(2)^r} +1\) We need to prove that last digit of this part is either 1/2/6/7 that part is even for sure so therefore we just gotta prove that its last digit is either 2 or 6 :/

OpenStudy (empty):

either \(a=b\) or \(a>b\) so those are our two cases to check. Mod 3 arithmetic we have: \[0^2=0\]\[1^2=1\]\[2^2=1\] so we know: \[3^a+3^b+1 = (3n+\alpha)^2\] where \(\alpha = 1\) or \(\alpha = 2\). Work it out a little bit and we get: \[3^a+3^b+1 = 9n^2+6n \alpha + \alpha^2\] This leads to two cases after we plug in our two choices of alpha, case 1: \[3^a+3^b+1 = 9n^2+6n + 1\]\[3^{a-1}+3^{b-1} = n(3n+2)\] Since \(a \ge b\) we must have: \(n=3^{b-1}\) \[3^{a-b}+1 = 3n+2 \] Contradiction and also not true mod 3 anyways. Next case! case 2: \[3^a+3^b+1 = 9n^2+12n + 4\]\[3^{a-1}+3^{b-1} =(3n+1)(n+1)\] Since 3n+1 is a factor then that must mean a>b and \(n+1 = 3^{b-1}\) This leads to a contradiction: \[3^{a-b}+1 = 3n+1 \]\[3^{a-b-1} =n \] Gross but works.

OpenStudy (rsadhvika):



OpenStudy (rsadhvika):

@Empty how \(a\ge b\) implies \(n = 3^{b-1}\) ?

OpenStudy (rsadhvika):

@imqwerty that is a clever trick, but it doesn't work here see http://www.wolframalpha.com/input/?i=Table%5B3%5Ea%2B3%2B1,+%7Ba,+1,+10%7D%5D the expression 3^a + 3^b + 1 evaluates to numbers ending in all digits

OpenStudy (ikram002p):

take mod 4 ... only 4n and 4n+1 can be perfect squares 3^a+3^b+1 mod 4 = (-1)^a+(-1)^b+1 mod 4 cases a even b even 1+1+1 mod 4 can't be perfect square a odd b odd -1-1+1 mod 4 =-1 mod 4 can't be perfect square a odd b even (or b odd a even) -1+1+1=1 mod 4 need to check this case :-\

OpenStudy (ikram002p):

i'll try mod 12 only 0,1,4,9 perfect 4 and 0 are even so not considering them. 3^a+3^b+1 mod 3 =1 mod 3 3^a+3^b+1 mod 4 see we would have same problem with case a odd and b even :-\

OpenStudy (rsadhvika):

Very nice ! mod 8 will do that case too i think

OpenStudy (ikram002p):

lets find another way...

OpenStudy (rsadhvika):

First we observe that \(3^a+3^b + 1\) is odd

OpenStudy (rsadhvika):

Recall that an odd perfect square is congruent to 1 in mod 8

OpenStudy (rsadhvika):

So it is sufficient if we could establish below : \[3^a+3^b+1\not\equiv 1 \pmod{8}\]

OpenStudy (ikram002p):

hahaha damn so by contradiction assuming 3^a+3^b+1 = 1 mod 8 then 3^a+3^b| 8 brilliant !

OpenStudy (rsadhvika):

Easy to see that \(3^n \equiv 1,3\pmod{8}\)

OpenStudy (rsadhvika):

None of the combinations \(\{1,3\}\) add up to \(0 \pmod{8}\). So we're done !

OpenStudy (ikram002p):

cool!